OpenStudy (anonymous):

find cos x if sin^2 x - 1/cos x = -1

OpenStudy (whpalmer4):

is that supposed to be sin^2x - 1 --------- = -1 cos x ?

OpenStudy (anonymous):

yessir

OpenStudy (whpalmer4):

would have been better to write it as (sin^2 x - 1)/cos x = -1 what you wrote is actually sin^2 x - (1/cos x) = -1 by the order of operator precedence (multiplication and division happen before addition and subtraction)

OpenStudy (anonymous):

sorry about that! i didn't bother to use parenthesis because i didn't realize that it'd be different without them.

OpenStudy (whpalmer4):

If sin^2 x + cos^2 x = 1, then sin^2 x + cos^2 x - 1 = 0 sin^2 x - 1 = -cos^2 x does that give you any ideas?

OpenStudy (anonymous):

unfortunately not.

OpenStudy (whpalmer4):

you must be getting tired, too :-) sin^2 x - 1 = -cos^2 x your problem is (sin^2 x - 1) / cos x = -1 so we can write that as (-cos^2 x) / cos x = -1 now we can cancel common factors on the left, leaving us with -cos x = -1 cos x = 1 can you solve that?

OpenStudy (whpalmer4):

remember, cos^2 x is just cos x * cos x so we had -cos x * cos x ------------ = -1 cos x

OpenStudy (whpalmer4):

I do so wish that OpenStudy would fix the blinketyblank equation formatting!

OpenStudy (anonymous):

cosine of 0 is 1, so the answer is 0?

OpenStudy (whpalmer4):

let's try it out! we think x = 0 might be the answer. let's plug it into the original and see if it works sin x = sin 0 = 0 sin^2 0 = 0*0 = 0 sin^2 0 - 1 = 0 - 1 = -1 cos 0 = 1 -1/cos 0 = -1 / 1 = -1 -1 = -1 looks like it is a valid solution!

OpenStudy (anonymous):

perfect! thank you so much for your help tonight.

OpenStudy (whpalmer4):

you betcha! good night...